mhaythamrajab:

The circular bull’s-eye of an archery target has a diameter of 24 cm, which is surrounded by 4 concentric rings with a width of 12 cm each. draw the target in a coordinate plane with center at the origin. write the equations of the circles that form the boundaries of the different regions of the target.

Vocaloid:

Since all the circles will have their center at the origin, the equation for each circle is simply x^2 + y^2 = r^2 where r is the radius. so for the center bulls-eye, that has diameter 24, its radius is 24/2 = 12, making the equation x^2 + y^2 = 12^2 repeat this logic with the other circles